# Self-Distilled Masked Auto-Encoders are Efficient Video Anomaly Detectors

### v1

@@ 4.1 Experimental Setup
Data sets. We verify the performance of our method on three popular data sets for video anomaly detection: Avenue [47], ShanghaiTech [49] and UCSD Ped2 [51]. The training videos of these data sets contain only normal events, and the test ones include both normal and abnormal scenarios. ShanghaiTech is the largest one, with 270K frames for training and about 50K for testing. Avenue is the second largest, with about 15K frames for training and another 15K for testing. UCSD Ped2 holds a total of 4.5K frames, out of which 2.5K are used for training. To augment the training videos, we sample abnormal events from the UBnormal data set [1]. UBnormal is a synthetic (virtual) data set containing anomalies simulated by video game characters, which alleviates the burden of collecting anomalies from the real world. The probability of augmenting a frame is $0.25$.
Hyperparameters. The encoder module is formed of three CvT blocks, each with a projection size of 256 and four attention heads. The teacher decoder contains three CvT blocks, while the student decoder contains only one block. All decoder blocks have four attention heads and a projection dimension of 128. Since the data sets have different input resolutions and objects vary in size, we adapt the patch size to each data set. Thus, we set the patch size to $16\times 16$ on Avenue, $8\times 8$ on ShanghaiTech, and $4\times 4$ on UCSD Ped2. Regardless of the data set, the teacher network is trained for 100 epochs, while the student is trained for 40 epochs. We optimize the networks with Adam [39], using a learning rate of $10^{-4}$ and mini-batches of 100 samples.
Type Method Avenue Shanghai Ped2 FPS
@@
Object-centric [9] 91.6 92.5 83.8 90.5 - - -
[50] + [9] 91.6 92.4 83.6 90.6 - - -
[17] 86.4 - 71.6 - 97.8 - -
[25] 91.5 92.8 82.4 90.2 97.5 99.8 51
[26] 92.3 90.4 82.7 89.3 98.7 99.7 24
[65] + [26] 92.9 91.9 83.6 89.5 - - 31
[50] + [26] 93.2 91.8 83.3 89.3 - - -
[33] 87.4 90.4 78.7 84.9 94.3 97.8 -
[79] 92.2 - 84.3 - 99.0 - -
[89] 89.6 - 74.8 - 97.3 - -
Frame or cube level [5] 84.7 - 73.7 - 98.4 - -
[4] 87.1 - 75.9 - 96.5 - -
[27] 83.3 - 71.2 - 94.1 - 35
[35] 88.9 - - - - - -
[40] 90.0 - - - 96.6 - -
[45] 84.4 - - - 87.5 - -
[44] 85.1 81.7 72.8 80.6 95.4 - 28
[46] 89.9 93.5 74.2 83.2 99.3 - 12
[50] + [46] 89.5 93.6 75.2 83.8 - - -
[50] + [44] 89.1 84.8 74.6 83.3 - - -
[50] + [57] 86.4 86.3 70.6 80.3 - - -
[53] 86.9 - - - 96.2 - -
[57] 82.8 86.8 68.3 79.7 97.0 - 101
[56] 85.3 - 72.2 - 96.3 - 195
[58] 72.0 - - - 88.3 - -
[59] 87.2 - - - 93.0 - -
[63] - - - - 93.5 - -
[62] - - - - 88.4 - -
[65] + [44] 87.3 84.5 74.5 82.9 - - 26
[65] + [46] 90.9 92.2 75.5 83.7 - - 10
[65] + [57] 84.8 88.6 69.8 80.2 - - 95
[72] 84.6 - - - - - -
[73] - - - 76.5 - - -
[74] 89.6 - 74.7 - - - -
[76] 85.1 - 73.0 - 96.3 - -
[82] 87.0 - 79.3 - - - -
[84] - - 80.4 - - - -
[85] 86.6 - - - 96.9 - -
[90] 90.2 - - - 97.3 - -
[97] - - - - 91.0 - -
Ours 90.5 89.6 77.3 84.2 93.1 97.8 1670
Table 1: Micro and macro AUC scores (in %) of several state-of-the-art frame-level, cube-level and object-level methods versus our self-distilled masked AE on Avenue, ShanghaiTech and UCSD Ped2. The top three scores for each category of methods are shown in red, green, and blue. All reported running times are measured on a machine with an Nvidia GeForce GTX 3090 GPU with 24 GB of VRAM.
@@ 4.2 Results
Results on Avenue. Our method obtains a micro AUC score of $90.5\%$ on Avenue, being only $2.7\%$ below the state-of-the-art object-centric method. However, in the category of frame-level methods, we reach the second-best micro AUC score. Taking into account that our method is much faster than all other methods, we consider that its performance gap is more than acceptable. In Figure 4, we illustrate the anomaly scores for test video $04$ from Avenue. Here, our model is close to perfect, highlighting its ability to capture anomalies, such as people running.
@@
Figure 5: Predictions for test video $01\_0051$ from ShanghaiTech. The abnormal bounding boxes are given by the convex hull of the patches labeled as abnormal. Best viewed in color.
Results on ShanghaiTech. On ShanghaiTech, our method reaches the top macro AUC score and the third-best micro AUC score when compared with other frame-level frameworks. Object-centric methods generally surpass frame-level methods, but the former methods have much lower processing speeds (see Figure 2). In Figure 5, we illustrate the anomaly scores for video $01\_0051$ from the ShanghaiTech test set. Our model reaches an AUC of $97.03\%$ on this video, being able to flag and locate the abnormal event, namely riding a bike into a pedestrian area.
Figure 6: Predictions for video Test001 from UCSD Ped2. The abnormal bounding boxes are given by the convex hull of the patches labeled as abnormal. Best viewed in color.
Results on UCSD Ped2. Our framework obtains a micro AUC of $93.1\%$, being $6.2\%$ below the state-of-the-art performance of Liu et al. [46]. While being slightly below in terms of micro AUC, our macro AUC score is remarkably on par with the object-centric method of Ionescu et al. [33], and only $2\%$ below that of Georgescu et al. [25]. Considering that our FPS is more than $30$ times higher compared with these methods[25, 33] on the same GPU, our framework provides a clearly superior performance-speed trade-off. In Figure 6, we illustrate the anomaly scores for video Test001 from UCSD Ped2. Here, our model reaches an AUC of $100\%$, being able to perfectly differentiate between normal and abnormal events.
Ablation study. In Table 2, we illustrate the impact of each novel component on our model’s performance. The first model is a vanilla masked AE, which obtains a rather low performance on Avenue. Each and every component contributes towards boosting the performance of the vanilla model. We observe that self-distillation gives the highest boost in terms of the micro AUC. However, to surpass the $90\%$ milestone, it is mandatory to introduce the prediction of anomaly maps in the learning task. The self-distillation procedure gives us a few possible strategies to combine the outputs of the teacher and the student. Hence, we investigate this aspect and report the results in Table 3. We observe that the best micro AUC is obtained when we combine the teacher reconstruction error with the teacher-student discrepancy. The proportion of synthetic examples per mini-batch is another aspect that can influence our model’s performance. In Table 4, we report the micro and macro AUC scores for three possible augmentation levels. We observe that augmentation is always useful, but, for a better outcome, it requires a moderate percentage ($25\%$).
Motion-based Self- Synthetic Anomaly AUC
weights distillation data maps Micro Macro
84.0 85.6
84.8 86.3
88.5 86.0
88.5 86.9
90.5 89.6
Table 2: Impact of each novel component on the micro and macro AUC scores (in %), on the Avenue data set [47].
Strategy AUC
Micro Macro
Teacher 84.0 85.6
Teacher + Student 85.4 85.8
Teacher + Teacher-Student Difference 88.5 86.0
Teacher + Student + Teacher-Student Difference 86.9 85.8
Table 3: Impact of strategies to combine the outputs of the teacher and student models on the Avenue data set [47].
Measure Percentage of synthetic data
@@
Micro AUC 88.5 90.5 89.1 88.7
Macro AUC 86.0 89.6 88.1 86.3
Table 4: Impact of varying the proportion of synthetic anomalies on the Avenue data set [47].
@@
Performance-speed trade-off. In Figure 2, we compare our model with several other methods in terms of the performance-speed trade-off. This comparison undoubtedly shows that our method reaches a far better processing speed, while achieving fairly good performance. To strengthen this observation, we also compare the methods in terms of GFLOPs and number of parameters in Table 5. Our method proves to be significantly lighter.
@@ 5 Conclusion
In this work, we proposed a lightweight masked auto-encoder (3M parameters, 0.8 GFLOPs) for video anomaly detection, which learns to reconstruct tokens with high motion gradients. Our framework is based on self-distillation, leveraging the discrepancy between teacher and student decoders for anomaly detection. Moreover, we boost the performance of our model by introducing a data augmentation technique based on overlapping synthetic anomalies on normal training data. Our highly efficient framework reached an unprecedented speed of 1670 FPS, with a minimal performance gap with respect to the state-of-the-art object-centric approaches.
In future work, we aim to further close the accuracy gap between our model and state-of-the-art models by introducing more students and refining the training procedure.
Acknowledgments
Work supported by a grant of the Romanian Ministry of Education and Research, CNCS - UEFISCDI, project number PN-III-P2-2.1-PED-2021-0195, within PNCDI III.

### v2

@@ 4.1 Experimental Setup
Data sets. We verify the performance of our method on four data sets for video anomaly detection: Avenue [51], ShanghaiTech [53], UBnormal [1] and UCSD Ped2 [55]. ShanghaiTech is the largest data set, with 270K frames for training and about 50K for testing. UBnormal is the second largest, with about 116K training frames and 93K testing frames. Avenue is a popular benchmark containing 15K frames for training and another 15K for testing. UCSD Ped2 holds a total of 4.5K frames, out of which 2.5K are used for training. UBnormal is a benchmark that uses an open set evaluation, where training and test anomalies belong to disjoint category sets. For the other three data sets, the training videos contain only normal events, and the test ones include both normal and abnormal scenarios. To augment the normal training videos, we sample abnormal events from the UBnormal data set [1]. UBnormal is a synthetic (virtual) data set containing anomalies simulated by video game characters, which alleviates the burden of collecting anomalies from the real world. The probability of augmenting a frame from Avenue, ShanghaiTech or UCSD Ped2 is $0.25$.
Type Method Avenue Shanghai UBnormal Ped2 FPS
@@
Micro
Macro
Object-centric [10] 91.6 92.5 83.8 90.5 62.1 86.5 - - 20
[18] 86.4 - 71.6 - - - 97.8 - -
[26] 91.5 92.8 82.4 90.2 55.4 84.5 97.5 99.8 51
[27] 92.3 90.4 82.7 89.3 61.3 85.6 98.7 99.7 24
[33] - - 85.9 - 71.8 - - - -
[36] 87.4 90.4 78.7 84.9 - - 94.3 97.8 -
[40] 80.2 - 73.7 - 50.7 - - - -
[49] 89.9 93.5 74.2 83.2 - - 99.3 - 12
[50] 91.8 92.3 83.8 87.8 - - - - -
[54] + [10] 91.6 92.4 83.6 90.6 - - - - 20
[54] + [27] 93.2 91.8 83.3 89.3 - - - - 31
[54] + [49] 89.5 93.6 75.2 83.8 - - - - 10
[69] + [27] 92.9 91.9 83.6 89.5 - - - - 31
[69] + [49] 90.9 92.2 75.5 83.7 - - - - 10
[84] 92.2 - 84.3 - - - 99.0 - 35
[95] 89.6 - 74.8 - - - 97.3 - -
Frame or cube level [4] 87.1 - 75.9 - - - 96.5 - -
[5] 84.7 - 73.7 - - - 98.4 - -
[9] - - - - 68.5 80.3 - - 37
[28] 83.3 - 71.2 - - - 94.1 - 35
[37] 88.9 - - - - - - - -
[43] 90.0 - - - - - 96.6 - -
[47] 85.1 81.7 72.8 80.6 - - 95.4 - 28
[48] 84.4 - - - - - 87.5 - -
[54] + [47] 89.1 84.8 74.6 83.3 - - - - 26
[54] + [61] 86.4 86.3 70.6 80.3 - - - - 94
[57] 86.9 - - - - - 96.2 - -
[60] 85.3 - 72.2 - - - 96.3 - 195
[61] 82.8 86.8 68.3 79.7 - - 97.0 - 101
[62] 72.0 - - - - - 88.3 - -
[63] 87.2 - - - - - 93.0 - -
[66] - - - - - - 93.5 - -
[67] - - - - - - 88.4 - -
[69] + [47] 87.3 84.5 74.5 82.9 - - - - 26
[69] + [61] 84.8 88.6 69.8 80.2 - - - - 95
[76] 84.6 - - - - - - - -
[77] - - - 76.5 50.3 76.8 - - 56
[78] 89.6 - 74.7 - - - - - -
[80] 85.1 - 73.0 - - - 96.3 - -
[82] - - 76.1 - - - - - -
[87] 87.0 - 79.3 - - - - - -
[89] - - 80.4 - - - - - -
[90] 86.6 - - - - - 96.9 - -
[93] 90.1 - 78.6 - 62.7 - - - -
[96] 90.2 - - - - - 97.3 - -
[100] - - 78.9 - - - - - -
[103] - - - - - - 91.0 - -
Ours 91.3 90.9 79.1 84.7 58.5 81.4 95.4 98.4 1655
Table 1: Micro and macro AUC scores (in %) of several state-of-the-art frame-level, cube-level and object-level methods versus our self-distilled masked AE on Avenue, ShanghaiTech, UBnormal and UCSD Ped2. The top three scores for each category of methods are shown in red, green, and blue. All reported running times (including those of the baselines) are measured on a machine with an Nvidia GeForce GTX 3090 GPU with 24 GB of VRAM.
Hyperparameters. The encoder module is formed of three CvT blocks, each with a projection size of 256 and four attention heads. The teacher decoder contains three CvT blocks, while the student decoder contains only one block. All decoder blocks have four attention heads and a projection dimension of 128. Since the data sets have different input resolutions and objects vary in size, we adapt the patch size to each data set. Thus, we set the patch size to $16\times 16$ on Avenue, $8\times 8$ on ShanghaiTech and UBnormal, and $4\times 4$ on UCSD Ped2. Regardless of the data set, the teacher network is trained for 100 epochs, while the student is trained for 40 epochs. We optimize the networks with Adam [42], using a learning rate of $10^{-4}$ and mini-batches of 100 samples. The hyperparameters in Eq. (6) are set to $\alpha=0.4$, $\beta=0.3$ and $\gamma=0.3$, for all data sets.
@@ 4.2 Results
Results on Avenue. Our method obtains a micro AUC score of $91.3\%$ on Avenue, being only $1.9\%$ below the state-of-the-art object-centric method. Remarkably, in the category of frame-level methods, we reach the best micro and macro AUC scores. Taking into account that our method is much faster than all other methods, we consider that its performance is remarkable. In Figure 4, we illustrate the anomaly scores for test video $04$ from Avenue. Here, our model is close to perfect, highlighting its ability to capture anomalies, such as people running.
@@
Results on ShanghaiTech. On ShanghaiTech, our method reaches the top macro AUC score and the third-best micro AUC score, when compared with other frame-level frameworks. Object-centric methods generally surpass frame-level methods, but the former methods have much lower processing speeds (see Figure 2).
Results on UBnormal. In terms of the micro AUC, the best frame-level method on UBnormal is TimeSformer [9], which benefits from large-scale pretraining. Notably, our method obtains a higher macro AUC than TimeSformer. The micro AUC of our method is fairly close to the micro AUC levels of the better object-centric approaches. In terms of speed, our method is significantly faster than all the other methods reporting results on UBnormal.
Results on UCSD Ped2. Our framework obtains a micro AUC of $95.4\%$, being $3.9\%$ below the state-of-the-art performance of Liu et al. [49]. While being slightly below in terms of micro AUC, our macro AUC score is remarkably on par with the object-centric method of Ionescu et al. [36], and only $1.4\%$ below that of Georgescu et al. [26]. Considering that our FPS is more than $30$ times higher compared with these methods [26, 36] on the same GPU, our framework provides a clearly superior accuracy-speed trade-off.
Motion Self- Synthetic Anomaly Classif. Avenue Shanghai
weights distillation data maps head AUC
Micro Macro Micro Macro
84.0 85.6 69.7 80.1
84.8 86.3 71.3 80.9
88.5 86.0 76.3 83.8
88.5 86.9 77.0 83.0
90.5 89.6 77.3 84.2
91.3 90.9 79.1 84.7
Table 2: Impact of each novel component on the micro and macro AUC scores (in %), on Avenue [51] and ShanghaiTech [53].
Strategy Avenue Shanghai
AUC
Micro Macro Micro Macro
Teacher 84.0 85.6 74.8 82.3
Teacher + Student 85.4 85.8 75.1 82.1
Teacher + Teacher-Student Difference 88.5 86.0 76.3 83.8
Teacher + Student + Teacher-Student Difference 86.9 85.8 75.8 83.6
Table 3: Impact of strategies to combine the outputs of the teacher and student models on Avenue [51] and ShanghaiTech [53]. These results do not include the synthetic anomalies and the classification head.
Data set Measure Percentage of synthetic data
@@
Avenue Micro AUC 88.5 91.3 90.6 89.9
Macro AUC 86.0 90.9 89.4 87.7
Shanghai Micro AUC 76.3 79.1 77.9 77.7
Macro AUC 83.8 84.7 84.4 84.1
Table 4: Impact of varying the proportion of synthetic anomalies on the Avenue [51] and ShanghaiTech [53] data sets.
@@
Ablation study. In Table 2, we illustrate the impact of each novel component on our model’s performance. The first model is a vanilla masked AE, which obtains a rather low performance on both Avenue and ShanghaiTech. Each and every component contributes towards boosting the performance of the vanilla model. We observe that self-distillation gives the highest boost in terms of the micro AUC. However, to surpass the $90\%$ milestone on Avenue, it is mandatory to introduce the prediction of anomaly maps in the learning task. An additional boost is given by our classification head.
The self-distillation procedure gives us a few possible strategies to combine the outputs of the teacher and the student. Hence, we investigate this aspect and report the results in Table 3. We observe that the best micro AUC is obtained when we combine the teacher reconstruction error with the teacher-student discrepancy.
The proportion of synthetic examples per mini-batch is another aspect that can influence our model’s performance. In Table 4, we report the micro and macro AUC scores for three possible augmentation levels. We observe that augmentation is always useful, but, for a better outcome, it requires a moderate percentage ($25\%$).
Performance-speed trade-off. In Figure 2, we compare our model with several other methods in terms of the performance-speed trade-off. This comparison undoubtedly shows that our method reaches a far better processing speed, while achieving fairly good performance. To strengthen this observation, we also compare the methods in terms of GFLOPs and number of parameters in Table 5. We underline that the method of Gong et al. [28] might seem small in terms of the number of parameters, but it is slowed down by its input, which is formed of a cuboid constructed by stacking 16 consecutive frames. Moreover, the method relies on a memory module, where each memory slot records the features of one pixel in the activation maps. Although their method has twice as many parameters as our own, the large input volume and the sizable memory bank reduce the speed to 35 FPS. With an FPS of 1655, our method proves to be significantly lighter than all its competitors.
@@ 5 Conclusion
In this work, we proposed a lightweight masked auto-encoder (3M parameters, 0.8 GFLOPs) for video anomaly detection, which learns to reconstruct tokens with high motion gradients. Our framework is based on self-distillation, leveraging the discrepancy between teacher and student decoders for anomaly detection. Moreover, we boost the performance of our model by introducing a data augmentation technique based on overlapping synthetic anomalies on normal training data. Our highly efficient framework reached an unprecedented speed of 1655 FPS, with a minimal performance gap with respect to the state-of-the-art object-centric approaches.
Acknowledgments. Work supported by a grant of the Romanian Ministry of Education and Research, CNCS-UEFISCDI, project number PN-III-P2-2.1-PED-2021-0195, within PNCDI III.